(1 point) We consider the non-homogeneous problem y - y = -4 cos(x) First we consider the homogeneous problem y -y = 0 : =
Answer #1

Given differential equation is y-y = -4COSX D the auxillary equcobon is 2-Y = 0 2) the roots of quxillary equabon are 10,1 (- A (off – Bìnx + + Bình - Bosx = -( ( JY = cosx f-A-B] e sinx -6+A] = -4 cost = -A-B=4 A-B =0 solving this we get [B= 2] , [